This is a multi-part, activity coefficient problem.

Part 1: Calculate the solubility of Ba(IO3)2 in pure water. Ksp of Ba(IO3)2=1.57*10^-9.

Part 2: Calculate the solubility of the same Ba(IO3)2 in a 0.0025 M Al(NO3)3 solution.
Answer #1

pogt2; що Emoto, Кse o patic, - səx16°. Кер, Kse = (А1 (No2) Iss-хіб = (o-oos) (25) Чs2. («S7x109 O00ZS- s2 - 000coots? | S=